# Bridging Gaps, Building Futures: Advancing Software Developer Diversity and Inclusion Through Future-Oriented Research

### v1

@@ 4.4. Connections to Socioeconomic Understanding
Given the spread of IT across nearly all aspects of human life, we have to acknowledge that socioeconomics has an impact on our field. Diversity and inclusion are not merely SE problems. They are projections of much broader socioeconomic problems, which have been studied in research from multiple social science disciplines, to name a few, sociology, anthropology, education, etc. Compared to SE research, they have developed socioeconomic understandings of these issues at much larger scales (Babbie, 2020). Such large-scale understandings and small-scale context-focused research in SE could complement each other. By connecting with these societal scale understandings, we might better distinguish the unique problems in SE and common social problems, position ourselves and our research in the full social spectrum, understand diversity and inclusion problems’ socioeconomic roots, inspire novel interventions, and coordinate to tackle diversity and inclusion problems in SE as a part of global social forces. From a micro-perspective at the individual level, socioeconomic understandings could remind us to seek solutions for the workplace dystopia described in the S3.
Social scientists have established such societal scale understandings mostly by tracking the socioeconomic dynamics over relative long periods, represented by the major multi-wave, nationwide, or international surveys, such as the General Social Survey, [^10] https://gss.norc.org/ World Value Survey, [^11] https://www.worldvaluessurvey.org/ and the Bureau of Labor Statistics, [^12] https://www.bls.gov/cps/lfcharacteristics.htm to name a few. The results of these surveys could be integrated with our research through a number of different ways. The results could be used in quantitative analysis to identify potential relationships, e.g., the overall labor market dynamics and women’s involvement in SE, in which data from the Bureau of Labor Statistics could be used. They might also provide contextual information in qualitative inquiries, e.g., the World Value Survey may help SE researchers interpret qualitative data about the differences in women’s participation between the United States and China. Besides, when designing and delivering educational materials, these socioeconomic understandings can offer unique insights to help us better understand the audience.
@@
Benefits (Utopia) Harms (Dystopia) Research Directions
Better and more comprehensive understandings of diversity and inclusion issues within and beyond the software engineering industry Coordinated effort to address larger problems Potential impacts across traditional discipline boundaries
Incorrectly attributing SE-specific problems to general socioeconomic problems Taking a passive attitude to wait for socioeconomic changes Ignoring the research results from other disciplines Using socioeconomic factors as an excuse for the inaction
Exploring forms of connecting socioeconomic understandings with SE research in diversity and inclusion Developing customized interventions considering different groups’ socioeconomic backgrounds Coordinating with social scientists to address national/international diversity and inclusion issues Preparing for the potential socioeconomic changes resulting from recent progresses in generative AI
@@ 5.1. An Agenda of SDDI
Our four research themes and their research goals jointly provide an actionable agenda, bridging gaps in SDDI to make the two utopian scenarios possible and avoid the two dystopian scenarios. First, appropriate research on methodologies and metrics must be developed to effectively analyze various aspects of SDDI in SE. Specifically, two concrete aspects of SDDI that we believe are of particular importance are the intersectionality of software engineers and the connection of large-scale socioeconomic understandings to SE practice. These two aspects need to be studied and understood more thoroughly to suggest appropriate SDDI initiatives and actionable principles for SE practitioners. Finally, SE research connected to SDDI will not impact education or practice without successful knowledge transfer. Given the relevance of societal and governmental initiatives to SDDI, transferring to and from these groups requires dedicated focus. In summary, our research goals for the four themes are as follows:
@@
Research on SDDI typically deals with marginalized groups. As such, this type of research requires a constant focus on maximizing benefits while minimizing harms, especially to vulnerable groups. The contrast between utopia and dystopia in our scenarios highlights this fine balance in an extreme way. Metrics, diversity dimensions, and socioeconomic understanding can and have been used both to the benefit and to the disadvantage of various societal groups. As a research community, we must strike this balance in a responsible way.
@@ 5.2. The Impact of AI on SDDI
The recent progress in generative AI, exemplified by the advent of LLMs and the burgeoning trend of computing as a general education, holds transformative potential for SDDI research. These advances are poised to revolutionize every facet of our agenda, sparking new avenues of exploration and understanding across research, industry, and educational contexts to build a more inclusive future in software development. We should be mindful that AI-based solutions not only inherit traditional SDDI challenges, but also come with new ones. For example, solutions produced by GitHub Copilot for Chinese prompts were found to be subpar compared to their English and Japanese counterparts (Koyanagi et al., 2024). This might create obstacles for developers preferring to express themselves in Chinese. Similarly, using ChatGPT-like solutions often involves tinkering, which is known to be a more common learning strategy among men than among women (Beckwith et al., 2006; Burnett et al., 2016b). In contrast, ChatGPT-like solutions could help several groups of developers who might feel more comfortable posing their questions to a machine rather than asking people, e.g., neurodivergent developers who commonly face difficulties in communication (Morris et al., 2015).
@@
AI can also play a role in intersectionality, affecting SDDI in computing education and practice. In addition to the methodological benefits mentioned above, AI could play a positive role in SE education involving people with diverse identities. One of the major promises of AI is to provide personalized materials to people of different characteristics (Dong et al., 2023). Thus, AI could bring individualized learning experiences to individuals of certain intersectionality, such as gender-sensitive, accessible software development learning materials for students (see S2: Education Utopia). Moreover, conversational agents powered by AI techniques have the potential to create psychologically safe development environments in which individuals of certain intersectionality would not feel embarrassed when interacting with AI. However, researchers must keep in mind that individuals from different identities might interact with AI technologies differently or that they might benefit differently from AI. For instance, studies report developers who identify as female and LGBTQ+ have significantly lower intent to learn and adopt AI-assisted coding platforms, while software engineers of racial minorities have a higher intent to upskill but also more negative perceptions of AI compared to their counterparts (Pluralsight, 2023).
@@
Generative AI technologies would inevitably change the landscape of today’s socioeconomic situations. These technologies might create new disparities that have never been seen before. For example, well-paid professional labor markets such as software development might experience a reduction (Acemoglu et al., 2022). How can we make various minority group members suffer less if this happens? How can we reskill minority group members to participate in future work? How can we avoid further polarization of the labor market? To address these issues, SDDI researchers should collaborate with researchers from other disciplines to closely monitor the socioeconomic dynamics and develop forward-looking solutions. When it comes to educational contexts, AI could help to scale computing education to larger cohorts as the adoption of computing education rises across all disciplines (Association for Computing Machinery and IEEE Computer Society (IEEE-CS)(2020), ACM). In particular, AI can support learning for individuals who may not otherwise be able to receive such education due to their socioeconomic backgrounds—for example, learners from the Global South. However, access to advanced AI leads to a new digital divide (Khowaja et al., 2024). While LLMs are almost ubiquitous in high-income economies, reliable Internet access remains a big issue in many low-income economies (Kula et al., 2022), contributing to the Workplace Dystopia (Scenario S3) described in Section 3. In this process, people from minority groups may lose their human anchors, which gives them a sense of belonging that keeps them in the area (see Scenario S4: Education Dystopia). Hence, SDDI research should not underestimate the potential socioeconomic changes caused by fast-evolving generative AI technologies.
@@ 6. Conclusions
Software affects almost all areas of modern life, affecting user behavior, well-being, and decision-making. Software engineers design and develop software applications—yet the diversity of software development teams does not represent the diversity of the global population. To this end, this work presents insights from SE researchers and practitioners on challenges and research opportunities to promote software developer diversity and inclusion (SDDI). We provide motivating utopian and dystopian scenarios describing the effects of diversity on SE practice and education in 2030 and discuss ways to promote SDDI through research methodologies, intersectionality, knowledge transfer, and socioeconomic understanding to navigate the changing landscape of software development. We further briefly discuss the potential impact of the recent progress in generative AI and the burgeoning trend of computing as a general education.
7. Acknowledgements
Some initial ideas presented in this paper were originally developed at the 2023 Software Developer Diversity and Inclusion Workshop. We thank the NII Shonan meeting for hosting the workshop and all workshop participants who contributed to the discussions and our original report: Denae Ford, Emerson Murphy-Hill, Magaret-Anne (Peggy) Storey, Rafael Prikladnicki, Yi (Oliver) Wang, Alexander Serebrenik, Grischa Liebel, Bogdan Vasilescu, Kelly Blincoe, Raula Gaikovina Kula, Gema Rodriguez-Perez, Cornelia Connolly, Jocelyn Simmonds, Alannah Oleson, Sonja Hyrynsalmi, Stefan Stanciulescu, Chris Brown, Ayushi Rastogi, Bianca Trinkenreich, Ray Harvey, Sebastian Baltes, Meiyappan Nagappan, and Zaina Saadeddin. Love you aaaaaaaaalllll!

### v2

@@ 4.4. Connections to Socioeconomic Understanding
Given the spread of IT across nearly all aspects of human life, we have to acknowledge that socioeconomics has an impact on our field. Diversity and inclusion are not merely SE problems. They are projections of much broader socioeconomic problems, which have been studied in research from multiple social science disciplines, e.g., sociology, anthropology, and education. Compared to SE research, they have developed socioeconomic understandings of these issues at much larger scale (Babbie, 2020). Such large-scale understandings and small-scale context-focused research in SE could complement each other. By connecting with these societal scale understandings, we might better distinguish the unique problems in SE and common social problems, position ourselves and our research in the full social spectrum, understand diversity and inclusion problems’ socioeconomic roots, inspire novel interventions, and coordinate to tackle diversity and inclusion problems in SE as a part of global social forces. From a micro-perspective at the individual level, socioeconomic understandings could remind us to seek solutions for the undesirable workplace scenario described in Scenario S3.
Social scientists have established such societal scale understandings mostly by tracking the socioeconomic dynamics over relatively long periods, represented by the major multi-wave, nationwide, or international surveys, such as the General Social Survey [^14] https://gss.norc.org/, World Value Survey [^15] https://www.worldvaluessurvey.org/, and the Bureau of Labor Statistics [^16] https://www.bls.gov/cps/lfcharacteristics.htm. The results of these surveys could be integrated with our research through a number of different ways. The results could be used in quantitative analysis to identify potential relationships, e.g., the overall labour market dynamics and women’s involvement in SE, in which data from the Bureau of Labor Statistics could be used. They might also provide contextual information in qualitative inquiries, e.g., the World Value Survey may help SE researchers interpret qualitative data about the differences in women’s participation between the United States and China. Besides, when designing and delivering educational materials, these socioeconomic understandings can offer unique insights to help us better understand the audience.
@@
Benefits Harms Research Directions
Better and more comprehensive understandings of diversity and inclusion issues within and beyond the software engineering industry Coordinated effort to address larger problems Potential impacts across traditional discipline boundaries Incorrectly attributing SE-specific problems to general socioeconomic problems Taking a passive attitude to wait for socioeconomic changes Ignoring the research results from other disciplines Using socioeconomic factors as an excuse for the inaction Exploring forms of connecting socioeconomic understandings with SE research in diversity and inclusion Developing customized interventions considering different groups’ socioeconomic backgrounds Coordinating with social scientists to address national/international diversity and inclusion issues Preparing for the potential socioeconomic changes resulting from recent progresses in generative AI
@@ 5.1. An Agenda of SDDI
Our four research themes and their research goals jointly provide an actionable agenda, bridging gaps in SDDI to make the two desirable “utopian” scenarios possible and avoid the two undesirable “dystopian” scenarios. First, appropriate research on methodologies and metrics must be developed to effectively analyze various aspects of SDDI in SE. Specifically, two concrete aspects of SDDI that we believe are of particular importance are the intersectionality of software engineers and the connection of large-scale socioeconomic understandings to SE practice. These two aspects need to be studied and understood more thoroughly to suggest appropriate SDDI initiatives and actionable principles for SE practitioners. Finally, SE research connected to SDDI will not impact education or practice without successful knowledge transfer. Given the relevance of societal and governmental initiatives to SDDI, transferring to and from these groups requires dedicated focus. In summary, our research goals for the four themes are as follows:
@@
Research on SDDI typically deals with marginalized groups. As such, this type of research requires a constant focus on maximizing benefits while minimizing harms, especially to vulnerable groups. The contrast between the desired and the undesired state in our scenarios highlights this fine balance in an extreme way. Metrics, diversity dimensions, and socioeconomic understanding can and have been used both to the benefit and to the disadvantage of various societal groups. As a research community, we must strike this balance in a responsible way.
@@ 5.2. The Impact of AI on SDDI
Figure 4. SE 2030 Roadmap for SDDI
Although our main themes were methodologies and metrics, intersectionality, large-scale socioeconomic data, and knowledge transfer, two cross-cutting themes emerged from the group work at the original workshop, namely the two AI-related topics AI & SDDI and AI & CS education. Therefore, we add to our roadmap for 2030 for SDDI a particular discussion on AI (Figure 4).
Recent progress in generative AI, exemplified by the advent of LLMs and the burgeoning trend of computing as a general education, holds transformative potential for SDDI research. These advances are poised to revolutionize every facet of our agenda, sparking new avenues of exploration and understanding across research, industry, and educational contexts to build a more inclusive future in software development. We should be mindful that AI-based solutions not only inherit traditional SDDI challenges, but also come with new ones.
Languages: For example, solutions produced by GitHub Copilot for Chinese prompts were found to be subpar compared to their English and Japanese counterparts (Koyanagi et al., 2024). This might create obstacles for developers preferring to express themselves in Chinese. Moreover, while AI-generated contents or AI-based translation techniques lower the barriers for non-native English speakers’ participation, using they may also lead to potential biases towards those non-native speakers regarding their capabilities (Wang et al., 2024).
Cognitive Styles: Similarly, using ChatGPT-like solutions often involves playful experimentation known as “tinkering”, a learning strategy more common among men than among women (Beckwith et al., 2006; Burnett et al., 2016b). Interestingly enough, a recent study has found men believing that ChatGPT did not personalize their experience enough and did not allow global customization of its behaviour, which the authors attribute to the insufficient amount of tinkering supported by ChatGPT (Choudhuri et al., 2024). Tinkering is only one of the aspects of ChatGPT-like solutions where perceptions of individuals of different genders might diverge. For example, women tend to consider the decision-making process of ChatGPT to be less transparent than men (Choudhuri et al., 2024). In contrast, ChatGPT-like solutions could help several groups of developers who might feel more comfortable posing their questions to a machine rather than asking people, for example, neurodivergent developers who often face difficulties in communication (Morris et al., 2015).
Cultural Aspects: Moreover, Atari et al. have shown that LLM responses to psychological measures are not aligned with large-scale cross-cultural data (Atari et al., 2023). They have also shown that the performance of LLMs in cognitive psychological tasks is most similar to that of people from WEIRD societies.
Learning: In a study with more than 3,000 software engineers, Hicks et al. have found that the “AI Skill Threat”, that is the struggle with adapting to AI-assisted work, is more common for racially minoritized developers (Hicks et al., 2024). That group also rated the quality of the output of AI-assisted coding tools significantly lower than other groups (Hicks et al., 2024).
In summary, although aspects such as gender bias in software design (Huff, 2002) and gender difference in human-computer interaction (Czerwinski et al., 2002; Tan et al., 2003) and programming environments (Beckwith and Burnett, 2004) have been discussed already in the early 2000s, it took years, if not decades, for the broader topic of diversity, equity, and inclusion to reach the main discourse in the software engineering research community. With this paper, we want to raise awareness that, as motivated above, GenAI inherits traditional SDDI challenges and comes with new ones. We hope that both researchers and practitioners consider the existing body of knowledge on SDDI topics when developing and improving GenAI models and tools (Daniela Damian, Kelly Blincoe, Denae Ford, Alexander Serebrenik, Zainab Masood, 2024).
@@
AI can also play a role in intersectionality, affecting SDDI in computing education and practice. In addition to the methodological benefits mentioned above, AI could play a positive role in SE education involving people with diverse identities. One of the major promises of AI is to provide personalized materials to people of different characteristics (Dong et al., 2023). Thus, AI could bring individualized learning experiences to individuals of certain intersectionality, such as gender-sensitive, accessible software development learning materials for students (see S2: Desired Scenario in Education). Moreover, conversational agents powered by AI techniques have the potential to create psychologically safe development environments in which individuals of certain intersectionality would not feel embarrassed when interacting with AI. However, researchers must keep in mind that individuals from different identities might interact with AI technologies differently or that they might benefit differently from AI. For instance, studies report developers who identify as female and LGBTQ+ have significantly lower intent to learn and adopt AI-assisted coding platforms, while software engineers of racial minorities have a higher intent to upskill but also more negative perceptions of AI compared to their counterparts (Pluralsight, 2023).
@@
Generative AI technologies would inevitably change the landscape of today’s socioeconomic situations. These technologies might create new disparities that have never been seen before. For example, well-paid professional labor markets such as software development might experience a reduction (Acemoglu et al., 2022). How can we make various minority group members suffer less if this happens? How can we reskill minority group members to participate in future work? How can we avoid further polarization of the labor market? To address these issues, SDDI researchers should collaborate with researchers from other disciplines to closely monitor the socioeconomic dynamics and develop forward-looking solutions. When it comes to educational contexts, AI could help to scale computing education to larger cohorts as the adoption of computing education rises across all disciplines (Association for Computing Machinery and IEEE Computer Society (IEEE-CS)(2020), ACM). In particular, AI can support learning for individuals who may not otherwise be able to receive such education due to their socioeconomic backgrounds—for example, learners from the Global South. However, access to advanced AI leads to a new digital divide (Khowaja et al., 2024). While LLMs are almost ubiquitous in high-income economies, reliable Internet access remains a big issue in many low-income economies (Kula et al., 2022), contributing to the Undesirable Workplace Scenario (S3) described in Section 3. In this process, people from minority groups may lose their human anchors, which gives them a sense of belonging that keeps them in the area (see Undesirable Scenario in Education, S4). Hence, SDDI research should not underestimate the potential socioeconomic changes caused by fast-evolving generative AI technologies.
@@ 6. Conclusions
Software affects almost all areas of modern life, affecting user behaviour, well-being, and decision-making. Software engineers design and develop software applications—yet the diversity of software development teams does not represent the diversity of the global population. To this end, this work presents insights from SE researchers and practitioners on challenges and research opportunities to promote software developer diversity and inclusion (SDDI). We also offer a SE 2030 Roadmap for SDDI to encourage new research approaches for diversity and inclusion in software development. We provide motivating desired (“utopian”) and undesired (“dystopian”) scenarios describing the effects of diversity on SE practice and education in 2030 and discuss ways to promote SDDI through research methodologies, intersectionality, knowledge transfer, and socioeconomic understanding to navigate the changing landscape of software development. We further briefly discuss the potential impact of the recent progress in generative AI and the burgeoning trend of computing as a general education.
The initial themes presented in this paper were developed during a research meeting in 2023. We picked up these six themes in this paper, four of them discussed as dedicated research themes and the remaining two as a part of the undesirable/desirable scenarios and the discussion. These themes represent our collective opinions on the key challenges and research opportunities for promoting SDDI. The prioritization approach used through the 1-2-4-ALL structure, combined with the extensive collective experience of the FSE SE 2030 workshop participants, builds confidence that these themes are indeed the most relevant ones for future research endeavours. Furthermore, our discussion adds relevant existing research to support the importance of the themes. However, ultimately an argument could be made for a different focus and/or for different research themes.
Acknowledgements
Some initial ideas presented in this paper were originally developed at the 2023 Software Developer Diversity and Inclusion Workshop. We thank the NII Shonan meeting for hosting the workshop and all workshop participants who contributed to the discussions and our original report. Rafael Prikladnicki thanks CNPq in Brazil.